OpenStudy (anonymous):

Solve for x. |2x + 6| - 4 = 20

OpenStudy (justnick09):

9=x

OpenStudy (anonymous):

x=9,x=-15

OpenStudy (amorfide):

First of all make the absolute part of the equation, the subject. Then you know that the the absolute part can be the positive or negative value of what the equation is equal to, so from that you can make two equations, and solve for x

OpenStudy (amorfide):

for example, if I have |x+3|=6 x+3 = 6 x+3=-6 x=3, or x=-9